OpenStudy (anonymous):

If a silo contains 340L of milk. A small hose drains the milk out at a rate of 40L per minute. A second hose used to fill the silo with milk at a rate of 110L in a minute and a half. How long will it take to to have 700L of milk in the silo? How could it be expressed as an equation?

OpenStudy (matt101):

First let's find the net movement of milk. We know the output is 40 L/min. We're given an input, but over 1.5 mins. We need to know what it is in 1 min to be able to compare it to the output. We can do this by finding the ratio: x/1 = 110/1.5 --> x = 73.3 L/min. This means the net flow will be 73.3 L/min - 40 L/min = 33.3 L/min INTO the silo. We want to find how long it will take to get to 700 L, but we already have 340 L in the silo. This means what we really want to find is how long it will take to get to 700 L - 340 L = 360 L. We can find the time from t = 340 L / 33.3 L/min = 10.2 min. The general equation for this would be: V = 340 + 33.3t, which means that your final volume (V) will be 340 L plus 33.3 L/min times the time (t) you leave the silo to fill. Does that make sense? If you have any questions please let me know!

OpenStudy (anonymous):

Isn't t meant to be equal to 360/33.3 ? Thx for your help :)